OpenStudy (anonymous):

13. [8 −3] [-4 6] how do i find the inverse with steps pls :)

OpenStudy (anonymous):

to inverse 2x2: 1. matrix in form [a b] [c d] 2. to inverse you have to do the following \[(1/ad-bc)*\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] that will be the inverse

OpenStudy (anonymous):

ok thank u for that but waht is the answer becuase i have been trying that i just dont understand how to pug in

OpenStudy (anonymous):

haha \[\left[\begin{matrix}.16667 & .0833333 \\ .111111& .22222222\end{matrix}\right]\]

OpenStudy (anonymous):

\[\left[\begin{matrix}1/6 & 1/12 \\ 1/9 & 2/9\end{matrix}\right]\]